OpenStudy (anonymous):

use a half angle identity to find each exact value. cos 195

OpenStudy (austinl):

Hello there! Welcome to OpenStudy! Quick question about your question, is this in radians or degrees? I am assuming degrees.

OpenStudy (akashdeepdeb):

cos 195 = cos 390/2 we know cos 390 = cos (360 + 30) = cos 30 = sqrt 3/2 Now we know that \(2cos^2 \theta = 1 + cos 2 \theta\) From Half angle Identity \(2cos^2 \frac{\theta}{2} = 1 + cos \theta\) Therefore we can write the answer as \(2cos^2 \frac{390}{2} = 1 + cos ~390\) OR, \(cos^2 \frac{390}{2} = \frac{2 + \sqrt{3}}{4}\) OR, \(cos \frac{390}{2} = cos~ 195 ~~= ~\sqrt{\frac{2 + \sqrt{3}}{4}}\) Understood? :)

OpenStudy (anonymous):

thanks i was close but not close enough

OpenStudy (anonymous):

:)

OpenStudy (akashdeepdeb):

:)

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY !! :)

OpenStudy (anonymous):

thank you so much, i'm glad to be here :P

OpenStudy (akashdeepdeb):

:D AWESOME!